Find f'(x) if

f(x)= √[In(e2x + e-2x)]

f(x) = e-2x sin2x



1
Expert's answer
2021-09-09T07:54:54-0400

(1) Given, "f(x)= \sqrt{ln(e^{2x}+e^{-2x})}"


Differentiating with respect to x,

"f'=\frac{df}{dx}=\frac{1}{2\sqrt{ln(e^{2x}+e^{-2x})}}\frac{(2e^{2x}-2e^{-2x})}{(e^{2x}+e^{-2x})} = \frac{(e^{2x}-e^{-2x})}{(e^{2x}+e^{-2x})\sqrt{ln(e^{2x}+e^{-2x})}}"



(2) Given, "f(x) = e^{-2x} sin2x"


Differentiating both sides with respect to x,


"f'(x) = \frac{df}{dx}=2e^{-2x}cos2x -2e^{-2x}sin2x = 2[e^{-2x}cos2x-e^{-2x}sin2x]"